Previous

Calendar

November

2005

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

30

JD303

31

JD304(windy)
Halloween (boo!)

1

JD305 (blustery)
Site Visit (SpB): Details , Photos (25)

at 11:00 mst, rebooted russter2

at ~15:30 MST, swapped span gas (old cylinder=CC114714 (S-M) , new cylinder= CC156367)

at 14:35 MST, Laser dewar topped off (to 35.5 cm)

at 15:00 MST, setup 3 TCs connected to tc23x to branch

Compact Rental

2

JD306
Site Visit (AT): Details

ozone flux instrument shutdown (and taken off the tower)

slow ozone instrument (o3.slow) turned off

at 12:00 MST, hi-rate data archiving stopped on russter2 (restarted at 15:55 MST)

3

JD307
Lab Meeting at 12:15 (TR)

4

JD308
CME Meeting at FLAB (8:30-10:30)

5

JD309
Site Visit (KR): Details

laser dewar topped off

east/west robot: 4-flask swapped (KR)

6

JD310

7

JD311

8

JD312(windy)
Site Visit (SpB, LSD, MW): Details , Photos (16)

soils sampled. (MW)

from 11:27-11:44 MST, Changed profiler (li-6251) dessicant

Jeep Rental

9

JD313

10

JD314
Lab Meeting at 12:15 (LL)

11

JD315

12

JD316 (snow)
Site Visit (SS, KR): Details

Sean Schaeffer replenishes laser cylinders.

13

JD317 (snow)

14

JD318 (snow)

15

JD319

16

JD320

17

JD321
Lab Meeting at 12:15 (NT)

18

JD322

19

JD323

20

JD324 (calm/clear)
Site Visit (SpB, KR, JM, J): Details , Photos (32)

at 13:00 MST, swiped kryton hygrometer

NOTE: alignment of 21.5m csat had changed slightly (due to high winds)

pfp installed in the shed (JM)

laser dewar topped off, 12cm to 25.5cm (SpB)

east/west robots: flasks swapped (KR)

Compact Rental

21

JD325

22

JD326
SpB Off

23

JD327

24

JD328
CU Holiday (TURKEY DAY)

25

JD329
Site Visit (KR): Details

laser dewar topped off

east/west robot: not finished, no swap (KR)

CU Holiday (day after TURKEY DAY)

SpB Returns

26

JD330

27

JD331

28

JD332

29

JD333

30

JD334(snow)

1

JD335(snow)
Lab Meeting at 12:15 (SpB)

2

JD336(snow)

3

JD337(snow)